# Transformers as Unsupervised Learning Algorithms: A study on Gaussian Mixtures

### v1

@@ 2.1 Preliminaries
2.2 The TGMM architecture
A recent line of work xie2021explanation ; garg2022can ; bai2023tfstats ; akyrek2023what ; li2023transformers has been studying the capability of transformer that functions as a data-driven algorithm under the context of in-context learning (ICL). However, in contrast to the setups therein where inputs consist of both features and labels, under the unsupervised GMM setup, there is no explicitly provided label information. Therefore, we formulate the learning problem as learning an estimation algorithm instead of learning a prediction algorithm as in the case of ICL. A notable property of GMM is that the structure of the estimand depends on an unknown parameter $K$, which is often treated as a hyper-parameter in GMM estimation titterington_85 ; mclachlan2000finite . For clarity of representation, we define an isotropic Gaussian mixture task as $\mathcal{T}=({\boldsymbol{\theta}},{\mathbf{X}},K)$, where ${\mathbf{X}}$ is a i.i.d. sample generated according to ground truth ${\boldsymbol{\theta}}$ according to the isotropic GMM law and $K$ is the configuration used during estimation which we assume to be the same as the number of components of the ground truth ${\boldsymbol{\theta}}$. The GMM task is solved via applying some algorithm $\mathcal{A}$ that takes ${\mathbf{X}}$ and $K$ as inputs and outputs an estimate of the ground truth $\widehat{{\boldsymbol{\theta}}}=\mathcal{A}({\mathbf{X}};K)$.
In this paper, we propose a transformer-based architecture, transformers-for-Gaussian-mixtures (TGMM), as a GMM task solver that allows flexibility in its outputs, while at the same time being parameter-efficient, as illustrated in Fig. 1: A TGMM model supports solving $s$ different GMM tasks with $K\in\mathcal{K}:={\{K_{1},\ldots,K_{s}\}}$. Given inputs $N$ data points ${\mathbf{X}}$ and a structure configuration of the estimand $K$. TGMM first augments the inputs with auxiliary configurations about $K$ via concatenating it with a task embedding ${\mathbf{P}}=\text{embed}(K)$, i.e., ${\mathbf{H}}=\left[{\mathbf{X}}||{\mathbf{P}}\right]$, and use a linear $\operatorname{Readin}$ layer to project project the augmented inputs onto a shared hidden representation space for several estimand structures ${\{K_{1},\ldots,K_{s}\}}$, which is then manipulated by a shared transformer backbone that produces task-aware hidden representations. The TGMM estimates are then decoded by task-specific $\operatorname{Readout}$ modules. More precisely, with target decoding parameters of $K$ components, the $\operatorname{Readout}$ module first perform an attentive-pooling operation lee2019set :
where ${\mathbf{V}}_{o},{\mathbf{K}}_{o}\in\mathbb{R}^{(d+K)\times D}$, ${\mathbf{Q}}_{o}\in\mathbb{R}^{(d+K)\times K}$. The estimates for mixture probability are then extracted by a row-wise mean-pooling of the first $K$ rows of $\mathbf{O}$, and the estimates for mean vectors are the last $d$ rows of $\mathbf{O}$. We wrap the above procedure as ${\{\widehat{\pi}_{k},\widehat{\mu}_{k}\}}_{i\in[K]}=\operatorname{Readout}_{{\boldsymbol{\Theta}}_{\tt out}}{\left({\mathbf{H}}\right)}$. TGMM is parameter-efficient in the sense that it only introduces extra parameter complexities of the order $O(sdD)$ in addition to the backbone. We give a more detailed explanation of the parameter efficiency of TGMM in appendix Appendix B. We wrap the TGMM model into the following form:
2.3 Meta training procedure
We adopt the meta-training framework as in garg2022can ; bai2023tfstats and utilize diverse synthetic tasks to learn the TGMM model. In particular, during each step of the learning process, we first use a TaskSampler routine (described in Algorithm 1) to generate a batch of $n$ tasks, with each task having a probably distinct sample size. The TGMM model outputs estimates for each task, i.e., ${\{\widehat{\mu}_{k},\widehat{\pi}_{k}\}}_{k\in[K]}={\rm TGMM}_{{\boldsymbol{\Theta}}}{\left({\mathbf{X}};K\right)}$. Define $\widehat{{\boldsymbol{\pi}}}\mathrel{\mathop{:}}={\{\widehat{\pi}_{k}\}}_{k\in[K]}$ and $\widehat{{\boldsymbol{\mu}}}\mathrel{\mathop{:}}={\{\widehat{\mu}_{k}\}}_{k\in[K]}$. For a batch of tasks ${\{\mathcal{T}_{i}\}}_{i\in[n]}={\{{\mathbf{X}}_{i},{\boldsymbol{\theta}}_{i},K_{i}\}}_{i\in[n]}$, denote by ${\boldsymbol{\theta}}_{i}={\boldsymbol{\mu}}_{i}\cup{\boldsymbol{\pi}}_{i}$ and $\widehat{{\boldsymbol{\theta}}}_{i}=\widehat{{\boldsymbol{\mu}}}_{i}\cup\widehat{{\boldsymbol{\pi}}}_{i}={\rm TGMM}_{{\boldsymbol{\Theta}}}{\left({\mathbf{X}}_{i};K_{i}\right)}$, $i\in[n]$. Then the learning objective is thus:
where $\ell_{\mu}$ and $\ell_{\pi}$ are loss functions for estimation of $\mu$ and $\pi$, respectively. We will by default use square loss for $\ell_{\mu}$ and cross entropy loss for $\ell_{\pi}$. Note that the task sampling procedure relies on several sampling distributions $p_{\mu},p_{\pi},p_{N},p_{K}$, which are themselves dependent upon some global configurations such as the dimension $d$ as well as the task types $\mathcal{K}$. We will omit those dependencies on global configurations when they are clear from context. The (meta) training procedure is detailed in Algorithm 2.
Algorithm 1 TaskSampler
1:sampling distributions $p_{\mu},p_{\pi},p_{N},p_{K}$.
3:Sample a GMM task according to the type of task
[…]
4:Sample the size of inputs $N\sim p_{N}$.
[…]
Algorithm 2 (Meta) Training procedure for TGMM
1:task dimension $d$, task types $\mathcal{K}=\{K_{1},\ldots,K_{s}\}$, number of tasks per step $n$, number of steps $T$.
2:Initialize a TGMM model ${\rm TGMM}_{{\boldsymbol{\Theta}}^{(0)}}$.
3:for $t=1:T$ do
4:     Sample $n$ tasks ${\{\mathcal{T}_{i}\}}_{i\in[n]}$ independently using the TaskSampler routine 1.
7:end for
8:return Trained model ${\rm TGMM}_{{\boldsymbol{\Theta}}^{(T)}}$.
3 Experiments
3.1 Experimental Setup
We pick the default backbone of TGMM similar to that in garg2022can ; bai2023tfstats , with a GPT-2 type transformer encoder radford2019language of $12$ layers, $4$ heads, and $128$-dimensional hidden state size. The task embedding dimension is fixed at $128$. Across all the experiments, we use AdamW loshchilov2017fixing as the optimizer and use both learning rate and weight decay coefficient set to $10^{-4}$ without further tuning. During each meta-training step, we fix the batch size to be $64$ and train $10^{6}$ steps. For the construction of TaskSampler, the sampling distributions are defined as follows: For $p_{K}$, We sample $K$ uniformly from ${\{2,3,4,5\}}$; For $p_{\mu}$, given dimension $d$ and number of components $K$, we sample each component uniformly from $[-5,5]^{d}$. Additionally, to prevent collapsed component means ndaoud2022sharp , we filter the generated mean vectors with a maximum pairwise cosine similarity threshold of $0.8$; For $p_{\pi}$, given $K$, we sample each $\pi_{k}$ uniformly from $[0.2,0.8]$ and normalize them to be a probability vector; For $p_{N}$, Given a maximum sample size $N_{0}$, we sample $N$ uniformly from $[N_{0}/2,N_{0}]$. The default choice of $N_{0}$ is $128$. During evaluation, we separately evaluate $4$ tasks with $2,3,4,5$ components, respectively. With a sample size of $128$ and averaging over $1280$ randomly sampled tasks.
where $\tilde{\sigma}$ is the permutation such that $\tilde{\sigma}=\arg\min_{\sigma}\sum_{k\in[K]}\left\|{\widehat{\mu}_{{\sigma}(i)}-\mu_{i}}\right\|^{2}$. We obtain the permutation via solving a linear assignment program using the Jonker-Volgenant algorithm crouse2016implementing . We also report all the experimental results under two alternative metrics: cluster-classification accuracy and log-likelihood in Section E.2.
3.2 Results and findings
RQ1: Effectiveness
We compare the performance of a learned TGMM with the classical EM algorithm and spectral algorithm under $4$ scenarios where the problem dimension ranges over ${\{2,8,32,128\}}$. The results are reported in Fig. 2. We observe that all three algorithms perform competitively (reaching almost zero estimation error) when $K=2$. However, as the estimation problem gets more challenging as $K$ increases, the EM algorithm gets trapped in local minima and underperforms both spectral and TGMM. Moreover, while the spectral algorithm performs comparably with TGMM, it cannot handle cases when $K>d$, which is effectively mitigated by TGMM, with corresponding performances surpassing those of the EM algorithm. This demonstrates the effectiveness of TGMM for learning an estimation algorithm that efficiently solves GMM problems.
Shifts in sample size $N$ Under this scenario, we evaluate the learned TGMM model on tasks with sample size $N^{\text{test}}$ that are unseen during training.
Shifts in sampling distributions Under this scenario, we test the learned TGMM model on tasks that are sampled from different sampling distributions that are used during training. Specifically, we use the same training sampling configuration as stated in Section 3.1 and test on the following perturbed sampling scheme, with $\tilde{\mu}_{k}=\mu_{k}+\sigma_{p}\varepsilon_{k}$, where $\mu_{k}\overset{i.i.d.}{\sim}\operatorname{Unif}{\left([-5,5]^{d}\right)}$, $\varepsilon_{k}\overset{i.i.d.}{\sim}{\mathcal{N}}{\left(0,I_{d}\right)}$, $k\in[K]$ and ${\{\varepsilon_{k}\}}_{k\in[K]}$ is independent with ${\{\mu_{k}\}}_{k\in[K]}$.
Figure 3: Assessments of TGMM under test-time task distribution shifts I: A line with $N_{0}^{\text{train}}\rightarrow N^{\text{test}}$ draws the performance of a TGMM model trained over tasks with sample size randomly sampled in $[N_{0}^{\text{train}}/2,N_{0}^{\text{train}}]$ and evaluated over tasks with sample size $N^{\text{test}}$. We can view the configuration $128\rightarrow 128$ as an in-distribution test and the rest as out-of-distribution tests.
In Fig. 3, we report the assessments regarding shifts in sample size, where we set $N_{\text{test}}$ to be $128$ and vary the training configuration $N_{0}$ to range over ${\left\{32,64,128\right\}}$, respectively. The results demonstrate graceful performance degradation of out-of-domain testing performance in comparison to the in-domain performance. To measure performance over shifted test-time sampling distributions, we vary the perturbation scale $\sigma_{p}\in\{0,1,\ldots,10\}$ with problem dimension fixed at $d=8$. The results are illustrated in Fig. 4 along with comparisons to EM and spectral baselines. As shown in the results, with the increase of the perturbation scale, the estimation problem gets much harder. Nevertheless, the learned TGMM can still outperform the EM algorithm when $K>2$. Both pieces of evidence suggest that our meta-training procedure indeed learns an algorithm instead of overfitting to some training distribution.
RQ3: Flexibility Finally, we initiate two studies that extend both the TGMM framework and the (meta) learning problem of solving isotropic GMMs. In our first study, we investigated alternative architectures for the TGMM backbone. Motivated by previous studies park2024can that demonstrate the in-context learning capability of linear attention models such as Mamba series gu2023mamba ; dao2024transformers . We test replacing the backbone of TGMM with a Mamba2 dao2024transformers model with its detailed specifications and experimental setups listed in Section E.1. The results are reported in Fig. 5, suggesting that while utilizing Mamba2 as the TGMM backbone still yields non-trivial estimation efficacy, it is in general inferior to transformer backbone under comparable model complexity.
In our second study, we adapted TGMM to be compatible with more sophisticated GMM tasks via relaxing the isotropic assumption. Specifically, we construct anisotropic GMM tasks via equipping it with another scale sampling mechanism $p_{\sigma}$, where for each task we sample $\sigma\sim\text{softplus}(\tilde{\sigma})$ with $\tilde{\sigma}$ being sampled uniformly from $[-1,1]^{d}$. We adjust the output structure of TGMM accordingly so that its outputs can be decoded into both estimates of both mean vectors, mixture probabilities, and scales, which are detailed in Section E.1. Note that the spectral algorithm does not directly apply to anisotropic setups, limiting its flexibility. Consequently, we compare TGMM with the EM approach and plot results in Fig. 6 with the $\ell_{2}$-error metric accommodating errors from scale estimation. The results demonstrate a similar trend as in evaluations in the isotropic case, showcasing TGMM as a versatile tool in GMM learning problems.
Additional experiments We postpone some further evaluations to Appendix E, where we present a complete report consisting of more metrics and conduct several ablations on the effects of backbone scales and sample sizes.
4 Theoretical understandings
4.1 Understanding TGMM
We investigate the expressive power of transformers-for-Gaussian-mixtures(TGMM) as demonstrated in Section 3. Our analysis presents two key findings that elucidate the transformer’s effectiveness for GMM estimation: 1. Transformer can approximate the EM algorithm; 2. Transformer can approximate the power iteration of cubic tensor.
Transformer can approximate the EM algorithm. We show that transformer can efficiently approximate the EM algorithm (Algorithm A.1; see Appendix A) and estimate the parameters of GMM. Moreover, we show that transformer with one backbone can handle tasks with different dimensions and components simultaneously. The formal statement appears in Appendix C due to space limitations.
Theorem 1 (Informal).
Transformer can approximate power iteration of cubic tensor. Since directly implementing the spectral algorithm with transformers proves prohibitively complex, we instead demonstrate that transformers can effectively approximate its core computational step–the power iteration for cubic tensors (Algorithm 1 in JMLR:v15:anandkumar14b ; see Appendix A). Specifically, we prove that a single-layer transformer can approximate the iteration step:
$v^{(j+1)}=T{\left(I,v^{(j)},v^{(j)}\right)},~{}j\in\mathbb{N},$ (4)
where $I$ denotes the identity matrix and $T$ represents the given cubic tensor. For technical tractability, we assume the attention layer employs a ReLU activation function. The formal statement appears in Appendix D due to space limitations.
Theorem 2 (Informal).
There exists a $2L$-layer transformer ${\rm TF}_{\boldsymbol{\Theta}}$ with ReLU activation such that for any $d\leq d_{0}$, $T\in\mathbb{R}^{d\times d\times d}$ and $v^{(0)}\in\mathbb{R}^{d}$, given suitable embeddings, ${\rm TF}_{\boldsymbol{\Theta}}$ implements $L$ steps of Eq. 4 exactly.
We give some discussion of the theorems in the following remarks.
Remark 1.
(1) Theorem 1 demonstrates that a transformer architecture can approximate the EM algorithm for GMM tasks with varying numbers of components using a single shared set of parameters (i.e., one backbone ${\boldsymbol{\Theta}}$). This finding supports the empirical effectiveness of TGMM (RQ1 in Section 3.2). Additionally, Theorem 2 establishes that transformers can approximate power iterations for third-order tensors across different dimensions, further corroborating the model’s ability to generalize across GMMs with varying component counts.
(2) Theorem 1 holds uniformly over sample sizes $N$ and sampling distributions under mild regularity conditions, aligning with the observed robustness of TGMM (RQ2 in Section 3.2).
Remark 2.
Different "readout" functions are also required to extract task-specific parameters in our theoretical analysis, aligning with the architectural design described in Section 2.2. For further discussion, refer to Remark C.3 in Section C.2.
4.2 Proof Ideas
Proof Idea of Theorem 1. We present a brief overview of the proof strategy for Theorem 1. Our approach combines three key components: (1) the convergence properties of the population-EM algorithm pmlr-v125-kwon20a , (2) concentration bounds between population and sample quantities (established via classical empirical process theory), and (3) a novel transformer architecture construction. The transformer design is specifically motivated by the weighting properties of the softmax activation function, which naturally aligns with the EM algorithm’s update structure. For intuitive understanding, Fig. 7 provides a graphical illustration of this construction. The full proof is in Appendix C.
Proof Idea of Theorem 2. To approximate Eq. 4, we perform a two-dimensional computation within a single-layer transformer. The key idea is to leverage the number of attention heads $M$ to handle one dimension while utilizing the $Q,K,V$ structure in the attention layer. Specifically, let $T=(T_{i,j,m})_{i,j,m\in[d]}$ and $v^{(j)}=(v_{i}^{(j)})_{i\in[d]}$. Then, Eq. 4 can be rewritten as
$v^{(j+1)}=\sum_{j,m\in[d]}v_{j}v_{l}T_{:,j,m},$
where $T_{:,j,m}=(T_{i,j,m})_{i\in[d]}\in\mathbb{R}^{d}$. This operation can be implemented using $d$ attention heads, where each head processes a dimension of size $d$ (Fig. 8). The complete construction and proof are provided in Appendix D.
Figure 8: Illustration of implementing Eq. 4 via a multi-head attention structure, where colored boxes denote corresponding implementation components. Here $\sigma$ denotes the ReLU function.
5 Conclusion and discussions
In this paper, we investigate the capabilities of transformers in GMM tasks from both theoretical and empirical perspectives. We discuss the limitations and potential future research directions here.
Limitations and future work directions. First, while our theoretical analysis focuses on the approximation ability of transformers, the optimization dynamics remain unexplored. Second, approximating the full spectral algorithm (Algorithm A.2; see Appendix A) presents a significant challenge, which we leave for future work. Third, our study is limited to the expressivity of transformers on classical GMM tasks; exploring their performance on other unsupervised learning tasks is an interesting direction that warrants further investigation.
Appendix
[…]
Organization of the Appendix. In Appendix A, we formally present the GMM algorithms referenced in Section 2. We discuss the parameter efficiency of TGMM in Appendix B. Rigorous statements and proofs of Theorem 1 and Theorem 2 are provided in Appendix C and Appendix D, respectively. Additional experimental details are included in Appendix E.
Appendix A Algorithm Details
We state the classical algorithms of GMM mention in Section 2 in this section.
Algorithm A.1 EM algorithm for GMM
1:$\{X_{i},i\in[N]\}$, ${\boldsymbol{\theta}}^{(0)}=\{\pi_{1}^{(0)},\mu_{1}^{(0)},\cdots\pi_{K}^{(0)},\mu_{K}^{(0)}\}$
2:$j\leftarrow 0$
3:while not converge do
6:     $j\leftarrow j+1$
7:end while
Algorithm A.2 Spectral Algorithm for GMM
1:$\{X_{i},i\in[N]\}$
2:Compute the empirical moments $\hat{M}_{2}$ and $\hat{M}_{3}$ by
$\hat{M}_{2}$ $=\frac{1}{N}\sum_{i\in[N]}X_{i}\otimes X_{i}-I_{d},~{}$
4:Do first $K$-th robust tensor decomposition (Algorithm 1 in [3], see Algorithm A.3) for $\tilde{M}_{3}=\hat{M}_{3}{\left(W,W,W\right)}$:
$\tilde{M}_{3}\approx\sum_{k\in[K]}\lambda_{k}v_{k}^{\otimes 3}$
Algorithm A.3 Robust Tensor Power Method
1:symmetric tensor $T\in\mathbb{R}^{d\times d\times d}$, number of iterations $L$, $N$.
2:the estimated eigenvector/eigenvalue pair; the deflated tensor.
3:for $\tau=1$ to $L$ do
5:     for $t=1$ to $N$ do
6:         Compute power iteration update:
8:     end for
9:end for
11:Do $N$ power iteration updates (line 5) starting from $v_{N}^{(\tau^{*})}$ to obtain $\hat{v}$.
12:Set $\hat{\lambda}:=\tilde{T}(\hat{v},\hat{v},\hat{v})$.
13:return the estimated eigenvector/eigenvalue pair $(\hat{v},\hat{\lambda})$; the deflated tensor $\tilde{T}-\hat{\lambda}\hat{v}^{\otimes 3}$.
Appendix B On the parameter efficiency of TGMM
@@
Appendix C Formal statement of Theorem 1 and proofs
For analytical tractability, we implement $\operatorname{Readin}$ as an identity transformation and define $\operatorname{Readout}$ to extract targeted matrix elements hence they are both fixed functions. Actually, we also need "Readout" functions to get the estimated parameters for different tasks, see Remark C.3. To theoretical convenience, we use the following norm of transformers, which differs slightly from the definition in [4].
@@ Appendix C Formal statement of Theorem 1 and proofs
C.1 Formal statement of Theorem 1
@@
${\mathbf{H}}=\begin{bmatrix}\overline{X}_{1}&\overline{X}_{2}&\dots&\overline{X}_{N}\\ {\mathbf{p}}_{1}&{\mathbf{p}}_{2}&\dots&{\mathbf{p}}_{N}\end{bmatrix}\in\mathbb{R}^{D\times N},~{}{\mathbf{p}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\theta}}}_{i}\\ {\bf r}_{i}\end{bmatrix},~{}\overline{{\boldsymbol{\theta}}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\pi}}}_{\log}\\ \overline{\mu}_{i\%K}\\ c_{i\%K}\\ {\mathbf{0}}_{3K_{0}}\end{bmatrix}\in\mathbb{R}^{d_{0}+4K_{0}+1},~{}{\bf r}_{i}=\begin{bmatrix}{\mathbf{0}}_{\tilde{D}}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix}\in\mathbb{R}^{D-(2d_{0}+3K_{0}+1)},$ (5)
@@
To guarantee convergence of the EM algorithm, we adopt the following assumption for the initialization parameters, consistent with the approach in [24].
@@
Now we propose the theorem that transformer can efficient approximate the EM Algorithm (Algorithm A.1), which is the formal version of Theorem 1.
Theorem C.1.
Fix $10<\delta,\beta<$ 1 and $11/2<a<$ 1. Suppose there exists a sufficiently large universal constant $C\geq 128$ for which assumption (A1) holds. If $N$ is suffcient large and $\varepsilon\leq 1/{\left(100K_{0}\right)}$ sufficient small such that
$\frac{\tilde{c}_{1}}{(1-a)\pi_{\min}}\sqrt{\frac{R_{\max}(R_{\max}\vee d)\log{\left(\frac{24K}{\delta}\right)}}{N}}+\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{4N}{\delta})}{d}}\right)\right)N\varepsilon<\frac{1}{2}\left(a-\frac{1}{2}\right),$
@@
$\epsilon(N,\varepsilon,\delta,a)\mathrel{\mathop{:}}=\frac{\tilde{c}_{3}}{\left(1-a\right)\pi_{\min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}N}{\delta})}{N}}+\tilde{c}_{4}{\left(\frac{1}{\pi_{\min}}+N{\left(R_{\max}+d+\sqrt{2d\log\left(\frac{4N}{\delta}\right)}\right)}\right)}\varepsilon<a(1-\beta),$
@@
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},M=O(1),\log B_{{\boldsymbol{\Theta}}}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
Notably, Eq. 6 holds for all tasks satisfying $d\leq d_{0}$ and $K\leq K_{0}$, where the parameters of transformer ${\boldsymbol{\Theta}}$ remains fixed across different tasks ${\mathcal{T}}$.
Remark C.1.
From Theorem C.1, if we take $\varepsilon=\tilde{O}{\left(N^{-3/2}d^{-1/2}\right)}$ and $L=O{\left(\log N\right)}$, then we have
@@
Remark C.2.
We give some explanations for the notations in Theorem C.1. Define
@@
where ${\{\tilde{\mu}_{i}^{(j)},\tilde{\pi}_{i}^{(j)}\}}_{i\in[K]}$ are the parameters obtained at the $j$-th iteration of the population-EM algorithm (see Section C.3 for details). In the convergence analysis of the population-EM algorithm [24], it is shown that after the first iteration, the parameters lie in a small neighborhood of the true parameters with high probability (i.e., $D_{1}^{\rm pEM}\leq a$ for some $11/2\leq a<$ 1). Furthermore, the authors prove that the algorithm achieves linear convergence (i.e., $D_{j+1}^{\rm pEM}\leq\beta D_{j}^{\rm pEM}$ for $j\in\mathbb{N}_{+}$ and some $10<\beta<$ 1) with high probability if $D_{1}^{\rm pEM}\leq a$ holds. Following their notations, here $a$ represents the radius of the neighborhood after the first iteration, while $\beta$ is the linear convergence rate parameter. Finally, $\varepsilon$ controls the approximation error of the transformer.
C.2 Construction of transformer architecture and formal version of Fig. 7
@@
In this section, we give the transformer architecture construction in Theorem C.1. We denote $w_{ij}=w_{j}(X_{i}),i\in[N],k\in[K]$ in this subsection for simplicity. Recall that we have assumed that $d\leq d_{0}$, $K\leq K_{0}$ and $N$ is divisible by $K$($N/K\in\mathbb{N}$). We first restate the encoding formulas in Eq. 5:
@@
${\mathbf{H}}=\begin{bmatrix}\overline{X}_{1}&\overline{X}_{2}&\dots&\overline{X}_{N}\\ \overline{{\boldsymbol{\theta}}}_{1}&\overline{{\boldsymbol{\theta}}}_{2}&\dots&\overline{{\boldsymbol{\theta}}}_{N}\\ {\mathbf{p}}_{1}&{\mathbf{p}}_{2}&\dots&{\mathbf{p}}_{N}\end{bmatrix}\in\mathbb{R}^{D\times N},~{}\overline{{\boldsymbol{\theta}}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\pi}}}_{\log}\\ \overline{\mu}_{i\%K}\\ c_{i\%K}\\ \overline{{\mathbf{w}}}_{i}\\ \overline{{\mathbf{w}}}_{i\log}\\ \overline{{\boldsymbol{\pi}}}\end{bmatrix}\in\mathbb{R}^{d_{0}+4K_{0}+1},~{}{\mathbf{p}}_{i}\mathrel{\mathop{:}}=\begin{bmatrix}{\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix}\in\mathbb{R}^{D-(2d_{0}+3K_{0}+1)},$ (7)

### v2

@@ Appendix C Full Notation of Network Architecture
Appendix D On the parameter efficiency of TGMM
@@
Appendix E Limitations and future work directions
First, while our theoretical analysis focuses on the approximation ability of transformers, the optimization dynamics remain unexplored. This is a common theoretical challenge in ICL literature; see Bai et al. (2023); Lin et al. (2024); Giannou et al. (2025). Second, approximating the full spectral algorithm (Algorithm˜B.2; see Appendix˜B) presents a significant challenge, which we leave for future work. Third, our study is limited to the expressivity of transformers on classical GMM tasks; exploring their performance on other unsupervised learning tasks is an interesting direction that warrants further investigation.
Appendix F Formal statement of Theorem˜1 and proofs
For analytical tractability, we implement $\operatorname{Readin}$ as an identity transformation and define $\operatorname{Readout}$ to extract targeted matrix elements hence they are both fixed functions. Actually, we also need "Readout" functions to get the estimated parameters for different tasks, see Remark˜F.3. To theoretical convenience, we use the following norm of transformers, which differs slightly from the definition in Bai et al. (2023).
@@ Appendix F Formal statement of Theorem˜1 and proofs
F.1 Formal statement of Theorem˜1
@@
${\mathbf{H}}=\begin{bmatrix}\overline{X}_{1}&\overline{X}_{2}&\dots&\overline{X}_{N}\\ {\mathbf{p}}_{1}&{\mathbf{p}}_{2}&\dots&{\mathbf{p}}_{N}\end{bmatrix}\in\mathbb{R}^{D\times N},~{\mathbf{p}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\theta}}}_{i}\\ {\bf r}_{i}\end{bmatrix}$, (4)
$\overline{{\boldsymbol{\theta}}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\pi}}}_{\log}\\ \overline{\mu}_{i\%K}\\ c_{i\%K}\\ {\mathbf{0}}_{3K_{0}}\end{bmatrix}\in\mathbb{R}^{d_{0}+4K_{0}+1},~{\bf r}_{i}=\begin{bmatrix}{\mathbf{0}}_{\tilde{D}}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix}\in\mathbb{R}^{D-(2d_{0}+3K_{0}+1)},$ (5)
@@
To guarantee convergence of the EM algorithm, we adopt the following assumption for the initialization parameters, consistent with the approach in Kwon and Caramanis (2020).
@@
Now we propose the theorem that transformer can efficient approximate the EM Algorithm (Algorithm˜B.1), which is the formal version of Theorem˜1.
Theorem F.1.
Fix $0<\delta,\beta<1$ and $1/2<a<1$. Suppose there exists a sufficiently large universal constant $C\geq 128$ for which assumption (A1) holds. If $N$ is suffcient large and $\varepsilon\leq 1/{\left(100K_{0}\right)}$ sufficient small such that
$\frac{\tilde{c}_{1}}{(1-a)\pi_{\min}}\sqrt{\frac{R_{\max}(R_{\max}\vee d)\log{\left(\frac{24K}{\delta}\right)}}{N}}$
$\qquad+\tilde{c}_{2}\left(R_{\max}+d\left(1+\sqrt{\frac{2\log(\frac{4N}{\delta})}{d}}\right)\right)N\varepsilon<\frac{1}{2}\left(a-\frac{1}{2}\right),$
@@
$\epsilon(N,\varepsilon,\delta,a)$ $\mathrel{\mathop{:}}=\frac{\tilde{c}_{3}}{\left(1-a\right)\pi_{\min}}\sqrt{\frac{Kd\log(\frac{\tilde{C}N}{\delta})}{N}}$
$\qquad+\tilde{c}_{4}{\left(\frac{1}{\pi_{\min}}+N{\left(R_{\max}+d+\sqrt{2d\log\left(\frac{4N}{\delta}\right)}\right)}\right)}\varepsilon$
$<a(1-\beta),$
@@
$D=O(d_{0}+K_{0}),D^{\prime}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\varepsilon^{-1}\right)},$
$M=O(1),$
$\log B_{{\boldsymbol{\Theta}}}\leq\tilde{O}{\left(K_{0}R_{\max}{\left(R_{\max}+d_{0}\right)}\right)}.$
Notably, Eq.˜6 holds for all tasks satisfying $d\leq d_{0}$ and $K\leq K_{0}$, where the parameters of transformer $\boldsymbol{\Theta}$} remains fixed across different tasks $\mathcal{T}$}.
Remark F.1.
From Theorem˜F.1, if we take $\varepsilon=\tilde{O}{\left(N^{-3/2}d^{-1/2}\right)}$ and $L=O{\left(\log N\right)}$, then we have
@@
Remark F.2.
We give some explanations for the notations in Theorem˜F.1. Define
@@
where ${\{\tilde{\mu}_{i}^{(j)},\tilde{\pi}_{i}^{(j)}\}}_{i\in[K]}$ are the parameters obtained at the $j$-th iteration of the population-EM algorithm (see Section˜F.3 for details). In the convergence analysis of the population-EM algorithm (Kwon and Caramanis, 2020), it is shown that after the first iteration, the parameters lie in a small neighborhood of the true parameters with high probability (i.e., $D_{1}^{\rm pEM}\leq a$ for some $1/2\leq a<1$). Furthermore, the authors prove that the algorithm achieves linear convergence (i.e., $D_{j+1}^{\rm pEM}\leq\beta D_{j}^{\rm pEM}$ for $j\in\mathbb{N}_{+}$ and some $0<\beta<1$) with high probability if $D_{1}^{\rm pEM}\leq a$ holds. Following their notations, here $a$ represents the radius of the neighborhood after the first iteration, while $\beta$ is the linear convergence rate parameter. Finally, $\varepsilon$ controls the approximation error of the transformer.
F.2 Construction of transformer architecture and formal version of Fig.˜7
@@
In this section, we give the transformer architecture construction in Theorem˜F.1. We denote $w_{ij}=w_{j}(X_{i}),i\in[N],k\in[K]$ in this subsection for simplicity. Recall that we have assumed that $d\leq d_{0}$, $K\leq K_{0}$ and $N$ is divisible by KK($N/K\in\mathbb{N}$). We first restate the encoding formulas in Eq.˜4:
@@
${\mathbf{H}}=\begin{bmatrix}\overline{X}_{1}&\overline{X}_{2}&\dots&\overline{X}_{N}\\ \overline{{\boldsymbol{\theta}}}_{1}&\overline{{\boldsymbol{\theta}}}_{2}&\dots&\overline{{\boldsymbol{\theta}}}_{N}\\ {\mathbf{p}}_{1}&{\mathbf{p}}_{2}&\dots&{\mathbf{p}}_{N}\end{bmatrix}\in\mathbb{R}^{D\times N},$
$\overline{{\boldsymbol{\theta}}}_{i}=\begin{bmatrix}\overline{{\boldsymbol{\pi}}}_{\log}\\ \overline{\mu}_{i\%K}\\ c_{i\%K}\\ \overline{{\mathbf{w}}}_{i}\\ \overline{{\mathbf{w}}}_{i\log}\\ \overline{{\boldsymbol{\pi}}}\end{bmatrix}\in\mathbb{R}^{d_{0}+4K_{0}+1},~{\mathbf{p}}_{i}\mathrel{\mathop{:}}=\begin{bmatrix}{\mathbf{0}}_{D-(2d_{0}+5K_{0}+2)}\\ 1\\ {\mathbf{e}}_{i\%K}\end{bmatrix}\in\mathbb{R}^{D-(2d_{0}+3K_{0}+1)},$